# OT-VP: Optimal Transport-guided Visual Prompting for Test-Time Adaptation

### v1

@@ 4.2 Results
Tables 1 and 2 summarize our experimental outcomes in single-source and multi-source settings, respectively. In the single-source scenario, the model is trained on one domain and then adapted to another. The average accuracy is calculated across all 12 domain pairings for each trial. In the multi-source setting, one domain is designated as the target while the remaining three serve as sources. Each reported value is an average, with standard deviations across three separate trials. These trials vary by weight initialization and data splits. Comprehensive results can be found in the Appendix B.
@@
OT-VP consistently outperforms SOTA TTA methods across all datasets in both settings. Demonstrated by the results in Tables 1 and 2, OT-VP remarkably boosts the ERM model’s performance in both single-source and multi-source settings. Notably, OT-VP secures substantial improvements, with enhancements of $9.0\%$, $4.6\%$, and $1.4\%$ in the single-source scenario, and $0.7\%$, $2.4\%$, and $1.5\%$ in the multi-source scenario across the respective datasets: PACS, VLSC, and OfficeHome, highlighting its potent effect on model accuracy.
@@
The contrast in performance between alternative approaches and our OT-VP is noteworthy. The observation that Tent-BN and Tent-C exhibit only nominal or no improvements, and sometimes even lead to declines in performance on ViT models, aligns with expectations considering the original design of Tent. Tent was initially conceived for models relying heavily on batch normalization layers for domain adaptation by adjusting these layers to minimize prediction entropy. However, ViTs operate on a different architectural principle, lacking batch normalization layers in their standard configuration. This fundamental discrepancy underscores the challenges of directly applying methodologies designed for CNNs to ViTs without modifications or considerations of architectural differences. Furthermore, T3A consistently improves performance over Tent in both experimental setups, though its gains are lower than those achieved by OT-VP.
@@
Moreover, while ERM on ViTs demonstrates significant potential in adapting to unseen target domains, many prior DG and TTA methods struggle to enhance performance [15, 49]. This highlights the superior adaptability and efficacy of OT-VP.
@@
Figure 4: Prediction entropy across Algorithms in Single-Source and Multi-Source settings on PACS. In both settings, OT-VP demonstrates a marked reduction in entropy, outperforming Tent-C and Tent-BN, which target entropy minimization directly.
@@ 4.3 Ablation Study
Analysis on the objective functions. To explore the efficacy of our approach, we experimented with directly minimizing entropy during the prompt learning process. Interestingly, this direct focus on entropy reduction did not yield improvements. Instead, we observed a monotonic decrease in accuracy as entropy was reduced. This suggests that while reducing entropy might intuitively seem beneficial, focusing solely on this aspect can lead to overconfidence in the model’s predictions, including those that are incorrect. Such over-trust manifests as high confidence in erroneous predictions, underlining a potential pitfall of learning universal visual prompts purely on entropy reduction. These findings validate the effectiveness of our chosen objective function, which not only achieves the desired outcome but also implicitly manages entropy without compromising the reliability of predictions.
[…]
Our results highlight two crucial insights: Firstly, incorporating label information into the learned prompts significantly enhances the alignment between the source and target domains, corroborating findings from prior research [1]. This label augmentation facilitates a deeper, more meaningful adaptation process. Secondly, we observed a performance decrement with the reduction of $\lambda$ values, as illustrated in Fig. 5. However, performance reaches a plateau and stabilizes when $\lambda$ is sufficiently large, approaching an effective infinity in relation to feature distance, indicating an optimal balance in the adaptation mechanism.
The role of $\lambda$ in the cost function is crucial—it modulates the extent to which transportation across different classes is permissible. A higher $\lambda$ effectively restricts cross-class transport, serving as a preventative measure against such occurrences. Notably, setting $\lambda$ to infinity aligns the OT distance with the Wasserstein distance, underlining the significance of this parameter. Our analysis suggests that a larger $\lambda$ facilitates a reduction in overfitting risk during prompt optimization with stochastic gradient descent (SGD), especially when utilizing pseudo labels. This is achieved by imposing a stricter barrier against cross-class transportation, thereby ensuring a more disciplined adaptation process.
@@
Figure 5: Influence of hyperparameter $\lambda$ on Single-Source performance. Accuracy declines with smaller $\lambda$ values but stabilizes when $\lambda$ is large. These trends reveal the pivotal role of $\lambda$ in preventing cross-class transport and its impact on overfitting, particularly when using pseudo labels during prompt optimization. Notably, OT-VP’s $\lambda$ is $10^{8}$.
Analysis on prompt length. We opted for a prompt length of 4, aligning with DoPrompt’s configuration. Our experiments across varying prompt lengths—2, 4, 6, and 8—revealed that OT-VP’s performance remained relatively stable for smaller lengths (2, 4, 6), as indicated in Fig. 6. However, a notable decline in performance was observed with a length of 8, attributed to the increased risk of overfitting, a consequence of the absence of target labels. The decision to fix the prompt length is driven by the necessity to pre-select all parameters prior to any engagement with the target data, ensuring a standardized approach in our experiments.
Analysis on computation time. Despite employing SGD to optimize prompt tokens, OT-VP remains computationally efficient. We optimize only 4 prompt tokens over 5 epochs, utilizing a 20% hold-out split from both source and target data. While the multi-source setting involves processing triple the data to compute source representations compared to the single-source setting, the time required for both is nearly identical. Specifically, the average time is 39.5 seconds for the multi-source and 38.7 seconds for the single-source setting on the PACS dataset on our hardware. Moreover, the computational time is slightly influenced by the size of the datasets but remains relatively quick. For instance, in the PACS dataset, domain S, which has more than double the data of domain P, requires more processing time—51.7 seconds for S versus 34.5 seconds for P in the multi-source setting. Full results for PACS can be found in Table 3 in Appendix. In conclusion, OT-VP can efficiently learn prompts in both single and multi-source settings without significant computational overhead.
Figure 6: Effect of Prompt Length on OT-VP Performance. Consistency in performance is maintained for shorter prompt lengths (2, 4, 6), but an extended length of 8 prompts leads to a performance drop, likely due to overfitting in the absence of target labels.
@@ 5 Conclusion
In this paper, we present OT-VP, a novel test-time adaptation approach that leverages visual prompt learning for effective test-time adaptation. OT-VP stands out by adapting without altering the pre-trained model and effectively aligning the source and target domains, offering a more practical solution than established methods. Our experiments reveal OT-VP’s strengths: consistently outperforming existing TTA methods for ViT and DG prompt learning, and reducing prediction entropy to increase model confidence. By optimizing universal prompts for the target domain, OT-VP simplifies the adaptation process, enhancing the applicability of deep learning models in real-world settings.
Supplementary Material
@@ Appendix B Full Results
In this section, we present the computation time for OT-VP on the PACS dataset. For the single-source setting, the average computation time is calculated across three different sources.
Setting A C P S Average
@@
We present the implementation details of all baseline methods. For the implementation of DoPrompt, we set the prompt length to 4, with the coefficient $\lambda$ explored over the set $\{0.1,1,10\}$. The $M$ parameter for T3A is chosen from $\{1,5,20,50,100,\mathrm{N/A}\}$, while the configuration for Tent is determined from combinations of $\{0.1,1.0,10.0\}$ and $\{1,3\}$.
@@
In this section, we present the comprehensive outcomes in Tables 1 and 2. The experiments were conducted using three different seeds {0, 1, 2} within the DomainBed framework.
Algorithm A C P S
ERM A - 64.5$\pm$0.9 98.9$\pm$0.2 56.4$\pm$2.0
C 83.9$\pm$3.3 - 89.6$\pm$0.5 69.2$\pm$2.1
P 74.2$\pm$1.4 44.4$\pm$1.2 - 34.1$\pm$4.3
S 50.8$\pm$3.6 58.4$\pm$6.0 49.5$\pm$4.4 -
DoPrompt A - 64.6$\pm$0.6 98.5$\pm$0.7 56.5$\pm$4.7
C 84.1$\pm$2.2 - 90.1$\pm$1.7 74.0$\pm$2.2
P 75.6$\pm$2.0 46.2$\pm$0.7 - 35.2$\pm$3.8
S 46.4$\pm$4.9 55.0$\pm$5.8 45.1$\pm$3.0 -
Tent-C A - 64.6$\pm$0.6 98.9$\pm$0.2 56.3$\pm$5.8
C 83.9$\pm$3.0 - 89.6$\pm$0.2 69.0$\pm$2.0
P 74.4$\pm$1.9 44.5$\pm$1.3 - 33.7$\pm$4.9
S 50.4$\pm$3.5 58.3$\pm$6.0 49.0$\pm$3.7 -
Tent-BN A - 71.9$\pm$0.9 98.9$\pm$0.5 66.6$\pm$3.7
C 84.9$\pm$2.1 - 91.3$\pm$1.7 71.7$\pm$1.0
P 78.0$\pm$1.7 56.0$\pm$3.6 - 41.8$\pm$2.5
S 56.3$\pm$4.0 62.9$\pm$3.1 47.5$\pm$2.9 -
T3A A - 70.2$\pm$2.1 98.6$\pm$0.5 67.9$\pm$4.5
C 86.3$\pm$1.7 - 94.4$\pm$1.4 71.1$\pm$1.4
P 80.2$\pm$1.6 53.9$\pm$3.9 - 35.9$\pm$2.4
S 69.0$\pm$4.0 69.9$\pm$5.9 56.9$\pm$2.5 -
OT-VP-B A - 76.7$\pm$2.9 98.3$\pm$0.1 66.8$\pm$2.8
C 84.4$\pm$0.8 - 92.2$\pm$2.2 69.8$\pm$1.2
P 77.8$\pm$0.4 56.8$\pm$6.9 - 63.9$\pm$2.1
S 44.7$\pm$3.1 58.1$\pm$3.6 40.4$\pm$3.3 -
OT-VP A - 81.8$\pm$0.3 99.0$\pm$0.3 72.2$\pm$2.9
C 84.4$\pm$2.2 - 92.6$\pm$1.6 69.5$\pm$0.8
P 80.4$\pm$1.7 64.3$\pm$4.2 - 67.2$\pm$2.6
S 56.0$\pm$3.9 64.6$\pm$4.4 50.5$\pm$1.0 -
@@
Algorithm A C P S Average
ERM 91.3$\pm$0.5 82.3$\pm$0.7 98.9$\pm$0.3 75.6$\pm$0.4 87.0
DoPrompt 91.4$\pm$0.5 81.8$\pm$0.4 99.5$\pm$0.4 77.1$\pm$0.5 87.5
Tent-C 91.6$\pm$0.7 82.7$\pm$1.3 98.9$\pm$0.4 75.7$\pm$0.5 87.2
Tent-BN 91.1$\pm$0.3 82.4$\pm$1.3 98.3$\pm$0.4 76.8$\pm$0.7 87.2
T3A 91.5$\pm$0.5 81.8$\pm$0.9 99.0$\pm$0.5 77.4$\pm$0.3 87.4
OT-VP-B 91.2$\pm$0.5 81.8$\pm$0.7 99.4$\pm$0.5 77.4$\pm$0.5 87.3
OT-VP 92.0$\pm$0.5 83.0$\pm$0.4 99.2$\pm$0.4 76.4$\pm$0.2 87.7
@@
Algorithm C L S V
ERM C - 50.7$\pm$1.2 47.9$\pm$2.3 47.0$\pm$2.1
L 62.9$\pm$1.4 - 55.8$\pm$1.6 63.1$\pm$2.2
S 67.5$\pm$1.5 59.9$\pm$1.3 - 67.7$\pm$2.4
V 96.5$\pm$0.7 66.1$\pm$2.1 80.3$\pm$1.5 -
DoPrompt C - 53.4$\pm$1.5 50.0$\pm$1.6 50.5$\pm$1.8
L 71.7$\pm$2.1 - 57.8$\pm$1.5 70.1$\pm$1.3
S 67.8$\pm$1.9 62.5$\pm$0.8 - 66.2$\pm$1.9
V 98.6$\pm$0.8 62.0$\pm$1.9 78.8$\pm$2.1 -
Tent-C C - 50.4$\pm$1.7 48.3$\pm$1.3 47.0$\pm$2.9
L 70.3$\pm$1.9 - 55.8$\pm$2.4 63.2$\pm$2.1
S 67.2$\pm$2.1 59.8$\pm$1.7 - 67.9$\pm$2.0
V 96.5$\pm$0.6 66.0$\pm$2.1 88.2$\pm$1.8 -
Tent-BN C - 71.9$\pm$2.6 98.9$\pm$2.1 66.6$\pm$1.8
L 84.9$\pm$2.8 - 91.3$\pm$3.0 71.7$\pm$2.1
S 78.0$\pm$1.9 56.0$\pm$0.7 - 41.8$\pm$1.8
V 56.3$\pm$2.2 62.9$\pm$1.9 47.5$\pm$1.3 -
T3A C - 51.8$\pm$1.6 57.0$\pm$1.8 57.3$\pm$1.2
L 83.6$\pm$1.8 - 62.7$\pm$1.2 64.3$\pm$2.1
S 71.1$\pm$1.9 60.5$\pm$1.7 - 67.4$\pm$1.8
V 97.3$\pm$0.6 66.8$\pm$1.9 80.3$\pm$1.4 -
OT-VP-B C - 55.7$\pm$2.4 50.0$\pm$2.1 47.0$\pm$2.6
L 73.1$\pm$1.8 - 56.4$\pm$1.8 60.7$\pm$2.1
S 67.1$\pm$2.3 60.8$\pm$1.5 - 67.3$\pm$1.9
V 96.8$\pm$0.6 68.4$\pm$1.9 79.1$\pm$1.2 -
OT-VP C - 59.9$\pm$1.2 51.3$\pm$2.3 48.9$\pm$0.8
L 90.8$\pm$0.6 - 56.3$\pm$1.2 63.8$\pm$0.9
S 69.6$\pm$1.9 64.2$\pm$0.8 - 68.8$\pm$1.2
V 96.8$\pm$0.6 69.3$\pm$0.9 80.8$\pm$1.5 -
@@
Algorithm C L S V Average
ERM 96.5$\pm$0.5 65.5$\pm$1.1 75.2$\pm$0.8 76.7$\pm$0.4 78.5
DoPrompt 98.2$\pm$0.8 67.8$\pm$0.8 75.3$\pm$0.7 79.9$\pm$0.5 80.3
Tent-C 97.7$\pm$0.5 65.2$\pm$1.0 75.3$\pm$0.7 76.9$\pm$0.4 78.8
Tent-BN 86.3$\pm$1.0 66.2$\pm$0.9 68.8$\pm$0.7 72.6$\pm$0.5 73.5
T3A 97.3$\pm$0.5 65.6$\pm$0.9 78.0$\pm$0.7 79.3$\pm$0.7 80.0
OT-VP-B 96.8$\pm$0.4 71.9$\pm$0.8 75.2$\pm$0.8 76.9$\pm$0.4 80.2
OT-VP 96.8$\pm$0.4 73.1$\pm$1.1 76.8$\pm$0.8 77.0$\pm$0.5 80.9
@@
Algorithm A C P R
ERM A - 54.3$\pm$0.5 71.4$\pm$1.6 77.0$\pm$0.4
C 67.4$\pm$0.7 - 70.0$\pm$1.2 73.2$\pm$0.4
P 62.9$\pm$2.3 47.8$\pm$0.1 - 78.9$\pm$0.7
R 70.3$\pm$0.7 49.2$\pm$1.8 78.5$\pm$0.5 -
DoPrompt A - 52.1$\pm$3.2 71.7$\pm$0.1 79.0$\pm$0.5
C 67.4$\pm$0.1 - 71.7$\pm$0.4 75.5$\pm$0.1
P 66.8$\pm$1.4 47.8$\pm$0.1 - 79.0$\pm$0.1
R 72.2$\pm$1.1 48.8$\pm$1.9 79.6$\pm$0.7 -
Tent-C A - 54.5$\pm$0.6 69.9$\pm$1.5 76.9$\pm$0.6
C 66.9$\pm$0.7 - 69.6$\pm$1.4 73.6$\pm$0.2
P 62.9$\pm$0.3 47.9$\pm$0.1 - 78.6$\pm$0.7
R 71.0$\pm$0.5 47.1$\pm$1.7 79.9$\pm$0.3 -
Tent-BN A - 56.7$\pm$0.7 70.5$\pm$0.8 77.5$\pm$0.8
C 67.9$\pm$0.6 - 70.4$\pm$1.4 72.9$\pm$0.4
P 65.4$\pm$1.2 48.6$\pm$0.3 - 79.1$\pm$0.3
R 72.8$\pm$0.5 49.5$\pm$1.2 79.9$\pm$0.3 -
T3A A - 55.1$\pm$0.7 71.2$\pm$1.2 76.6$\pm$0.7
C 67.9$\pm$0.7 - 71.5$\pm$1.2 74.8$\pm$0.3
P 67.1$\pm$0.8 48.7$\pm$0.2 - 80.3$\pm$0.5
R 72.9$\pm$0.5 49.8$\pm$1.5 80.9$\pm$0.4 -
OT-VP-B A - 55.1$\pm$0.8 70.5$\pm$0.8 75.0$\pm$0.5
C 65.8$\pm$0.4 - 69.4$\pm$0.8 73.1$\pm$0.2
P 64.6$\pm$1.8 49.1$\pm$0.1 - 77.4$\pm$0.4
R 71.1$\pm$0.9 52.1$\pm$1.8 79.6$\pm$0.6 -
OT-VP A - 55.0$\pm$0.7 71.4$\pm$1.2 76.9$\pm$0.7
C 67.6$\pm$0.7 - 70.1$\pm$0.9 73.6$\pm$0.8
P 68.7$\pm$0.9 49.7$\pm$0.6 - 79.9$\pm$0.5
R 71.3$\pm$0.5 52.2$\pm$1.3 80.8$\pm$0.4 -

### v2

@@ 4.2 Experiments on PACS, VLCS, and OfficeHome
Tables 2 and 3 summarize our experimental outcomes for PACS, VLCS and OfficeHome in single-source and multi-source settings, respectively. Each reported value is an average across three separate trials. These trials vary by weight initialization and data splits. Comprehensive results can be found in the Appendix B.
@@
OT-VP consistently outperforms SOTA TTA methods across all datasets in both settings. Demonstrated by the results in Tables 2 and 3, OT-VP remarkably boosts the ERM model’s performance in both single-source and multi-source settings. Notably, OT-VP shows substantial improvements, with enhancements of $9.0\%$, $4.6\%$, and $1.4\%$ in the single-source scenario, and $0.7\%$, $2.4\%$, and $1.5\%$ in the multi-source scenario across the respective datasets: PACS, VLSC, and OfficeHome, highlighting its potent effect on model accuracy.
@@
The contrast in performance between alternative approaches and our OT-VP is noteworthy. The observation that Tent-BN and Tent-C exhibit only nominal or no improvements, and sometimes even lead to decrease in performance on ViT models, aligns with the expectations considering the original design of Tent. Tent was initially conceived for models relying heavily on batch normalization layers for domain adaptation by adjusting these layers to minimize prediction entropy. However, ViTs operate on a different architectural principle, lacking batch normalization layers in their standard configuration. This fundamental discrepancy underscores the challenges of directly applying methodologies designed for CNNs to ViTs without modifications or considerations of architectural differences. Furthermore, T3A and DePT consistently improve performance over Tent in both experimental setups, though their gains are smaller than those achieved by OT-VP.
@@
4.3 Experiments on ImageNet-C
We conducted further evaluations of our method on ImageNet-C, specifically at the highest severity level of 5. Results are detailed in Table 1 (top). For DePT, we cite results directly from [11], while outcomes for other methods are derived from three experimental runs. DoPrompt was excluded from this analysis because it necessitates changes to the pre-training process, which is incompatible with publicly available pre-trained models.
OT-VP boosts accuracy across all corrupted domains. Although T3A demonstrates performance improvements on PACS, VLCS, and OfficeHome, its effectiveness on ImageNet-C is limited. We hypothesize that the use of 1000 pseudo-prototypes and reliance on distance metrics for predictions complicates adaptation under these conditions. In contrast, our method, OT-VP, significantly enhances accuracy on ImageNet-C, outperforming T3A across all 15 domains and surpassing DePT in 12 out of 15 domains. Notably, the average accuracy improvement across these domains is 1.9% over DePT, despite our method only requiring the addition of four prompt tokens to the first transformer layer, whereas DePT integrates 50 prompt tokens across four different layers and updates classifier parameters.
OT-VP outperforms SOTA in online setting. We tested our method in an online setting where the model adapts to sequentially incoming target data batches, each used only once for training. We adapted our method by initially training on the first one percent of batches (about 8) for 50 steps and then reducing training to just one step for subsequent batches. This approach assumes all batches originate from the same domain, allowing us to continually refine a universal prompt for all batches, thus speeding up computation. If this domain consistency assumption does not hold, prompts could be updated across all batches for 50 steps to accommodate domain variability. The results, presented at the bottom of Table 1, show that OT-VP achieves an average accuracy of 65.7%, which, while 1.3% lower than the offline version, is still 2.1% higher than the online SOTA, DePT, and even slightly surpasses DePT offline by 0.6%.
In summary, while ERM on ViTs demonstrates significant potential in adapting to unseen target domains, many prior DG and TTA methods struggle to enhance performance [16, 52]. Our approach demonstrates considerable improvement across all four datasets, maintaining the entire pre-trained model’s parameters while learning only a minimal number of prompt tokens. This showcases our method’s efficiency and effectiveness in enhancing model adaptability without extensive modifications.
@@ 4.4 Ablation Study
Analysis of the objective functions. To explore the efficacy of our approach, we experimented with directly minimizing entropy during the prompt learning process. Interestingly, this direct focus on entropy reduction did not yield improvements. Instead, we observed a monotonic decrease in accuracy as entropy was reduced. This suggests that while reducing entropy might intuitively seem beneficial, focusing solely on this aspect can lead to overconfidence in the model’s predictions, including those that are incorrect. Such over-trust manifests as high confidence in erroneous predictions, underlining a potential pitfall of learning universal visual prompts purely on entropy reduction. These findings validate the effectiveness of our chosen objective function, which not only achieves the desired outcome but also implicitly manages entropy without compromising the reliability of predictions (see in App. B.5).
@@
Figure 4: (a) Influence of hyperparameter $\lambda$ on ImageNet-C. Accuracy declines with smaller $\lambda$ values but stabilizes when $\lambda$ is large. These trends reveal the pivotal role of $\lambda$ in preventing cross-class transport and its impact on overfitting, particularly when using pseudo labels during prompt optimization. (b) Effect of Prompt Length on OT-VP Performance. The performance of OT-VP exhibits only minor variations across different numbers of prompts, showing a robust improvement.
Analysis on $\lambda$. In our further investigations, we delved into the impact of the hyperparameter $\lambda$ within Eq. 7 on ImageNet-C. Our experiments spanned a range of $\lambda$ values: $[0,1,10,10^{2},10^{3},10^{4},10^{5}]$, where $\lambda=0$ and $\lambda=10^{4}$ corresponds to the OT-VP-B and OT-VP, respectively.
Our results highlight two crucial insights: Firstly, incorporating label information into the learned prompts significantly enhances the alignment between the source and target domains, corroborating findings from prior research [1]. This label augmentation facilitates a deeper, more meaningful adaptation process. Secondly, we observed a performance decrement with the reduction of $\lambda$ values, as illustrated in Fig. 4(a). However, performance reaches a plateau and stabilizes when $\lambda$ is sufficiently large, approaching an effective infinity in relation to feature distance, indicating an optimal balance in the adaptation mechanism. $\lambda=10^{4}$ is used for all main experiments.
Analysis on prompt length. We selected a prompt length of four, in alignment with DoPrompt’s configuration. Our experiments, which explored various prompt lengths—2, 4, 6, 8, and 10—demonstrated that the performance of the OT-VP varies only slightly with different numbers of prompts $l$ except for $l=2$, as illustrated in Fig. 4(a). This exception likely arises because two prompt tokens are insufficient for adequately capturing domain differences. Consequently, we standardized the prompt length at $l=4$ for all primary experiments. This decision ensures that all parameters are predetermined prior to any interaction with the target data, thereby facilitating a uniform approach across our experimental processes.
Algo. #Parameters #Forwards #Backwards Relative Time Acc.
ERM 0 782 0 0.6 55.5
Tent-BN 38,400 782 782 1.0 59.2
T3A 0 782 0 1.1 56.9
DePT 807,168 2,346 400 6.2 65.1
OT-VP-B 3,072 782 100 1.0 64.9
OT-VP 3,072 782 100 1.1 67.0
Table 4: Comparative analysis of trainable parameters and computational complexity on ImageNet-C in the Offline Setting. Note that T3A is an optimization-free method requiring prototype updates and generally exhibits suboptimal performance. Our method demonstrates significant improvements. It achieves these advancements with substantially fewer trainable parameters and reduced need for backward propagations than DePT.
Analysis of Trainable Parameters and Computational Complexity. Despite using SGD for optimizing prompt tokens, our OT-VP method remains computationally and memory-efficient. We detail the number of trainable parameters and computation times for all algorithms in Table 4. The trainable parameters for OT-VP consist of only four prompt tokens. In contrast, DePT requires not only a greater number of prompt tokens (50) but also updates to the classifier parameters. Our method outperforms DePT by 1.9% while utilizing only 0.4% of the trainable parameters compared to DePT. Furthermore, DePT necessitates additional forward passes due to the processing of three augmented images by both the teacher and student models. In contrast, our method requires only a single forward pass for each image. Additionally, the reduced number of trainable parameters allows our method to stabilize with fewer backward propagations. Consequently, OT-VP can be efficiently applied in real-world scenarios, both in terms of computation and memory usage.
@@ 5 Conclusion
In this paper, we introduce OT-VP, a novel test-time adaptation approach that utilizes visual prompt learning for effective test-time adaptation. Distinctively, OT-VP adapts without modifying the pre-trained model and effectively aligns the source and target domains, thus providing a more practical alternative to conventional methods. During adaptation, the process requires only the addition of four prompt tokens to the input token sequence, which are then updated using the optimal transport distance between the target representation and a precomputed source representation. We conduct extensive evaluations of the proposed method on both stylistic datasets (PACS, VLCS, OfficeHome) and corrupted datasets (ImageNet-C), demonstrating that learning merely four prompts can achieve state-of-the-art performance across multiple settings. Additionally, we confirm that OT-VP is both memory- and computation-efficient.
@@ Appendix B Full Results
In this section, we present the computation time for OT-VP on the PACS dataset. We optimize only 4 prompt tokens over 5 epochs, utilizing a 20% hold-out split from both source and target data. While the multi-source setting involves processing triple the data to compute source representations compared to the single-source setting, the time required for both is nearly identical. Specifically, the average time is 39.5 seconds for the multi-source and 38.7 seconds for the single-source setting on the PACS dataset on our hardware. Moreover, the computational time is slightly influenced by the size of the datasets but remains relatively quick. For instance, in the PACS dataset, domain S, which has more than double the data of domain P, requires more processing time—51.7 seconds for S versus 34.5 seconds for P in the multi-source setting. Full results for PACS can be found in Table 5 in Appendix. In conclusion, OT-VP can efficiently learn prompts in both single and multi-source settings without significant computational overhead. For the single-source setting, the average computation time is calculated across three different sources.
Setting A C P S Avg
@@
In this section, we provide details for the three stylistic datasets and one corrupted dataset. PACS [23] is composed of four domains: Photos, Art, Cartoon, and Sketch, containing 9,991 images in 7 classes. VLCS [8] comprises four real-world photographic datasets: VOC2007, LabelMe, Caltech, and SUN09, containing 10,729 images in 5 classes. OfficeHome [48] consists of four domains: Art, Clipart, Product, Real, containing 15,588 images in 65 classes. ImageNet-C comprises corrupted images in 15 types of corruption. We use the highest level of corruption (i.e. severity 5).
@@
We present the implementation details of our experiments. Following [14], we partition the data from each domain into training and validation splits of $80\%$ and $20\%$, respectively, utilizing the larger split for training and the smaller one for model selection. Our training approach for ERM adheres to the hyperparameters specified by [52], incorporating a dropout rate of $0.1$ and a weight decay of $10^{-2}$. The learning rate is $5\times 10^{-6}$ for PACS and VLCS, and $10^{-5}$ for OfficeHome.
For all baseline methods except for DePT, we use their official implementation [^2] https://github.com/DequanWang/tent  [^3] https://github.com/matsuolab/T3A  [^4] https://github.com/zhengzangw/DoPrompt. For the implementation of DoPrompt, we set the prompt length to 4, with the coefficient $\lambda$ explored over the set $\{0.1,1,10\}$. The $M$ parameter for T3A is chosen from $\{1,5,20,50,100,\mathrm{N/A}\}$, while the configuration for Tent is determined from combinations of $\{0.1,1.0,10.0\}$ and $\{1,3\}$. For DePT, we implement DePT-Group with $M=4$ stages and 50 prompts, adhering to the same hyperparameters specified for ImageNet-C in [11].
In the single-source scenario, the model is trained on one domain and then adapted to another. The average accuracy is calculated across all 12 domain pairings for each trial. In the multi-source setting, one domain is designated as the target while the remaining three serve as sources.
B.4
In this section, we present the comprehensive outcomes in Tables 2 and 3. The experiments were conducted using three different seeds {0, 1, 2} within the DomainBed framework. Tables 6, 8, and 10 display average results from three rounds for each source-target pair on PACS, VLCS, and OfficeHome, respectively, in the single-source setting. Similarly, Tables 7, 9, and 11 show the average results across three rounds for each target domain on PACS, VLCS, and OfficeHome in the multi-source setting.
Figure 5: Prediction entropy across TTA Algorithms in Single-Source and Multi-Source settings on PACS. In both settings, OT-VP demonstrates a marked reduction in entropy, outperforming Tent-C and Tent-BN, which target entropy minimization directly.
B.5
Algo. A C P S
ERM A - 64.5 98.9 56.4
C 83.9 - 89.6 69.2
P 74.2 44.4 - 34.1
S 50.8 58.4 49.5 -
DoPrompt A - 64.6 98.5 56.5
C 84.1 - 90.1 74.0
P 75.6 46.2 - 35.2
S 46.4 55.0 45.1 -
Tent-C A - 64.6 98.9 56.3
C 83.9 - 89.6 69.0
P 74.4 44.5 - 33.7
S 50.4 58.3 49.0 -
Tent-BN A - 71.9 98.9 66.6
C 84.9 - 91.3 71.7
P 78.0 56.0 - 41.8
S 56.3 62.9 47.5 -
T3A A - 70.2 98.6 67.9
C 86.3 - 94.4 71.1
P 80.2 53.9 - 35.9
S 69.0 69.9 56.9 -
DePT A - 70.1 98.4 64.4
C 83.9 - 90.1 69.3
P 76.6 49.7 - 36.1
S 51.4 63.4 52.3 -
OT-VP-B A - 76.7 98.3 66.8
C 84.4 - 92.2 69.8
P 77.8 56.8 - 63.9
S 44.7 58.1 40.4 -
OT-VP A - 81.8 99.0 72.2
C 84.4 - 92.6 69.5
P 80.4 64.3 - 67.2
S 56.0 64.6 50.5 -
@@
Algo. A C P S Gain
ERM 91.3 82.3 98.9 75.6 87.0
DoPrompt 91.4 81.8 99.5 77.1 87.5
Tent-C 91.6 82.7 98.9 75.7 87.2
Tent-BN 91.1 82.4 98.3 76.8 87.2
T3A 91.5 81.8 99.0 77.4 87.4
DePT 91.1 81.7 99.2 77.3 87.3
OT-VP-B 91.2 81.8 99.4 77.4 87.3
OT-VP 92.0 83.0 99.2 76.4 87.7
@@
Algo. C L S V
ERM C - 50.7 47.9 47.0
L 62.9 - 55.8 63.1
S 67.5 59.9 - 67.7
V 96.5 66.1 80.3 -
DoPrompt C - 53.4 50.0 50.5
L 71.7 - 57.8 70.1
S 67.8 62.5 - 66.2
V 98.6 62.0 78.8 -
Tent-C C - 50.4 48.3 47.0
L 70.3 - 55.8 63.2
S 67.2 59.8 - 67.9
V 96.5 66.0 88.2 -
Tent-BN C - 38.3 46.9 52.7
L 50.0 - 42.7 49.9
S 60.9 62.3 - 69.3
V 85.9 66.0 77.5 -
T3A C - 51.8 52.1 54.3
L 83.6 - 62.7 64.3
S 71.1 60.5 - 67.4
V 97.3 66.8 80.3 -
DePT C - 54.6 50.8 48.5
L 78.4 - 56.4 64.1
S 68.4 61.2 - 67.6
V 96.7 67.1 80.2 -
OT-VP-B C - 55.7 50.0 47.0
L 73.1 - 56.4 60.7
S 67.1 60.8 - 67.3
V 96.8 68.4 79.1 -
OT-VP C - 59.9 51.3 48.9
L 90.8 - 56.3 63.8
S 69.6 64.2 - 68.8
V 96.8 69.3 80.8 -
@@
Algo. C L S V Gain
ERM 96.5 65.5 75.2 76.7 78.5
DoPrompt 98.2 67.8 75.3 79.9 80.3
Tent-C 97.7 65.2 75.3 76.9 78.8
Tent-BN 86.3 66.2 68.8 72.6 73.5
T3A 97.3 65.6 78.0 79.3 80.0
DePT 96.6 69.2 76.7 77.8 80.1
OT-VP-B 96.8 71.9 75.2 76.9 80.2
OT-VP 96.8 73.1 76.8 77.0 80.9
@@
Algo. A C P R
ERM A - 54.3 71.4 77.0
C 67.4 - 70.0 73.2
P 62.9 47.8 - 78.9
R 70.3 49.2 78.5 -
DoPrompt A - 52.1 71.7 79.0
C 67.4 - 71.7 75.5
P 66.8 47.8 - 79.0
R 72.2 48.8 79.6 -
Tent-C A - 54.5 69.9 76.9
C 66.9 - 69.6 73.6
P 62.9 47.9 - 78.6
R 71.0 47.1 79.9 -
Tent-BN A - 56.7 70.5 77.5
C 67.9 - 70.4 72.9
P 65.4 48.6 - 79.1
R 72.8 49.5 79.9 -
T3A A - 55.1 71.2 76.6
C 67.9 - 71.5 74.8
P 67.1 48.7 - 80.3
R 72.9 49.8 80.9 -
DePT A - 54.9 70.2 75.9
C 68.1 - 70.2 73.5
P 64.7 48.6 - 79.2
R 71.0 49.6 79.1 -
OT-VP-B A - 55.1 70.5 75.0
C 65.8 - 69.4 73.1
P 64.6 49.1 - 77.4
R 71.1 52.1 79.6 -
OT-VP A - 55.0 71.4 76.9
C 67.6 - 70.1 73.6
P 68.7 49.7 - 79.9
R 71.3 52.2 80.8 -